OpenStudy (anonymous):

What is the slope of the line that passes through the points (–9, 5) and (–7, 10)? A. B. C. D.

OpenStudy (tennis5518):

WHat? the choice

OpenStudy (anonymous):

A. -5/2 B -2/5 C 2/5 D 5/2

OpenStudy (anonymous):

i know the slope is 5/2 but i dont know if it is negative

pooja195 (pooja195):

\[ \huge~\rm~\bf\color{blue}{ \frac{y_2-y_1}{x_2-x_1}=slope} \] \[ \huge~\rm~\bf\color{blue}{ \frac{10-5}{-7--9}=slope} \] \[ \huge~\rm~\bf\color{blue}{ \frac{10-5}{-7+9}=slope} \] \[ \huge~\rm~\bf\color{blue}{ \frac{10-5=5}{-7+9=2}=slope} \] It would remain positive -7+9=2 so there would be no reason for the slope to be negative

OpenStudy (tennis5518):

So basically there is no reason to put negative

OpenStudy (anonymous):

ok ty